OpenStudy (bluejay):

(2+x)/(x-4)+2=3/(2x-8). can someone explain? with work? thanks in advance.

OpenStudy (anonymous):

=(2+x)(2x-8)+2(x-4)(2x-8)=3(x-4) You can reach this by multiplying everything in the equation through by the denominator terms ie (x-4)(2x-8)

OpenStudy (bluejay):

i dont understand what you are trying to explain..